OpenStudy (anonymous):

How do you get the antiderivated of e^(2x). I'm having a massive smelly brain fart.

OpenStudy (anonymous):

derivative* (I can't even spell)

OpenStudy (amistre64):

lol ... or just multiply by 2/2

jimthompson5910 (jim_thompson5910):

hint: let u = 2x, so du = 2dx ----> dx = du/2 So now you need to integrate e^(u)(du/2)

OpenStudy (amistre64):

\[\frac{1}{2}\int 2e^{2x}dx\]

OpenStudy (amistre64):

what i do is think of where e^2x would have come from: well, e^2x of course, then derive it and adjust. [e^2x]' = 2 e^2x ; so we have to divide off a two to get our answer

OpenStudy (anonymous):

that is so much work. you know that it will look like \[?e^{2x}\] because the input will never change. so just think "what would i differentiate to get \[e^{2x}\] and you should come up with \[\frac{1}{2}e^{2x}\] by dint of reasoning

OpenStudy (amistre64):

any of this make any sense?

OpenStudy (anonymous):

I get the u-substitution version. Thanks guys